OpenStudy (anonymous):

The ratio of two complementary angles is 2:3, find the measures of each angle ?

OpenStudy (studygurl14):

Do you know what "complementary angles" means?

OpenStudy (anonymous):

i really not s:

OpenStudy (studygurl14):

If two angles are complementary, it means that they add up to 90°.

OpenStudy (studygurl14):

(refresh if you see diamonds)

OpenStudy (studygurl14):

Anyway, it says that two complementary angles have a ratio of 2:3. Do this: 2 +3 = 5 total parts. To find the measure of first angle: \(\large 90\Large\times\frac{2}{5}\) To find the measure of the second angle: \(\large 90\Large\times\frac{3}{5}\)

OpenStudy (studygurl14):

@Leyla_12 Do you understand?

OpenStudy (anonymous):

I'm confused why you put first\[\frac{ 2 }{ 5}\] and then 3/5 ?

OpenStudy (studygurl14):

Because you have two angles.

OpenStudy (studygurl14):

Remember you have the ratio of 2:3, meaning that if you split it into five parts, the first angle would make up two fifths of the total, and the second would make up three fifths of the total.

OpenStudy (studygurl14):

Does that make sense?

OpenStudy (anonymous):

So i have to multiply 90 x 2/5 and 90 x 3/5 crossed ?

OpenStudy (anonymous):

Nope. 90 * 2/5 and 90 * 3/5 will give you the angles right away. Think of it this way: You have 90 degrees to split into two angles so that their ratio is 2:3. It would make sense to divide 90 by 2+3=5, which becomes 1 fifth, or 1 part. Then give 2 of such parts to one angle and 3 of such parts to the second angle.